### v2

# Retrieve to Explain: Evidence-driven Predictions with Language Models

Abstract
Machine learning models, particularly language models, are notoriously difficult to introspect. Black-box models can mask both issues in model training and harmful biases. For human-in-the-loop processes, opaque predictions can drive lack of trust, limiting a model’s impact even when it performs effectively. To address these issues, we introduce Retrieve to Explain (R2E). R2E is a retrieval-based language model that prioritizes amongst a pre-defined set of possible answers to a research question based on the evidence in a document corpus, using Shapley values to identify the relative importance of pieces of evidence to the final prediction. R2E can adapt to new evidence without retraining, and incorporate structured data through templating into natural language. We assess on the use case of drug target identification from published scientific literature, where we show that the model outperforms an industry-standard genetics-based approach on predicting clinical trial outcomes.
Retrieve to Explain, R2E, Machine learning, Explainability, Retrieval, Language model, Shapley values, Human-in-the-loop, Multimodal, Drug discovery, Target identification, Genetics, Biomedical
1 Introduction
Recent research has shown that language models can act as knowledge bases, supplying answers to factual user queries on-demand using only the learned parameters (Petroni et al., 2019; Brayne et al., 2022). Language models can also be supplied with searchable knowledge bases for a retrieval-augmented approach to question answering (Chen et al., 2017; Lewis et al., 2020; Izacard & Grave, 2021).
Figure 1: R2E drug target identification example. R2E makes predictions based on retrieved evidence and provides explanations in terms of the evidence. Query: User queries are phrased in cloze-style, where [MASK] can be filled from a set of answers (named entities). Retrieval: R2E retrieves the evidence most relevant to the query for each possible answer. Prediction: The model generates a score for each answer based on the evidence. Explanation: Predictions are explained using Shapley values of the evidence. Here, the best evidence is indirect, based on the role of CD6 in mechanisms central to rheumatoid arthritis pathology.
Beyond answering factual queries, a searchable knowledge base could provide evidence for queries without known answers, including scientific research questions (e.g. What are some promising drug targets to treat rheumatoid arthritis?). By proposing new hypotheses supported by both direct and indirect scientific evidence, AI models could support and accelerate the scientific discovery process (Paliwal et al., 2020; Aliper et al., 2023; Sourati & Evans, 2023).
When acting on model hypotheses is costly or risky, an explainable model can mitigate risk by allowing a human expert to inspect the evidence and reasoning behind each prediction before acting on it (i.e. a human-in-the-loop setup). Explainability can also help to identify model flaws or systemic biases, leading to improved performance and task alignment (Kulesza et al., 2015).
Here, we introduce Retrieve to Explain (R2E), an approach for explainable language model predictions (Figure 1). Given a cloze-style user query, R2E first retrieves the most relevant evidence from an evidence corpus, partitioned according to each possible answer; we consider a set of answers comprised of named entities. The model then scores each answer based on its supporting evidence to generate a fully ranked list. The R2E architecture is designed to ensure that potential answers are represented explicitly in terms of their supporting evidence. In particular, the feature space is comprised of the evidence itself, allowing the use of feature attribution methods to estimate the contribution of each piece of evidence to the final prediction; here, we use Shapley values (Shapley et al., 1953; Lundberg & Lee, 2017). In addition to clarifying the prediction to the user, we show that this evidence-oriented approach allows model predictions to be updated without retraining by modifying the corpus, such as updating it with new evidence. Since R2E can generate a score for every answer in the answer set, it is particularly applicable in human-in-the-loop scenarios where many potential hypotheses should be prioritized for user review. In addition, this approach allows users to investigate negative, low scoring predictions.
With half of drugs failing to show efficacy when tested in human populations (Wong et al., 2019), we focus here on drug target identification, a critical early stage in drug discovery where specific genes or proteins (targets) are selected as the focus for developing treatments. For this task, we train R2E to score protein-coding genes by relevance to a user query based on a scientific literature corpus. We then augment the corpus with genetic associations by templating them into natural language, allowing the model to use both evidence sources. We show how Shapley values on the evidence correlate with Large Language Model (LLM) relevance assessments, which similarly correlate with human experts. Notably, when used to predict clinical trial outcomes, R2E significantly outperforms genetics evidence, a widely recognised predictor in the pharmaceutical industry (Nelson et al., 2015; Trajanoska et al., 2023). R2E outperforms the genetics baseline even when supplied only with genetics evidence, suggesting that representing gene-disease associations in natural language improves generalization. Finally, we show that explainability enables the use of LLMs to audit prediction reasoning, further improving performance.
Alongside the clinical trial outcomes, we evaluate the model on two additional target identification benchmarks and make all three new benchmarks publicly available (Appendix A).
2 Related work
2.1 Language Models with Retrieval
Many language models have been proposed to leverage retrieved training examples at inference time (Khandelwal et al., 2019; Karpukhin et al., 2020; Guu et al., 2020; Lewis et al., 2020; Yogatama et al., 2021; Izacard & Grave, 2021; Borgeaud et al., 2022; Izacard et al., 2022). R2E differs from these existing approaches by (1) scoring all possible answers in a given answer set and (2) attributing each answer’s score to evidence passages using Shapley values. This approach follows from the application: R2E is designed for answering research questions that merit deep user engagement (e.g. identifying drug targets for a disease) as opposed to typical factual recall tasks (e.g. identifying a country’s capital city). In the former case, there is a premium on generating multiple possible answers and explanations, allowing a human user to investigate inference results in depth.
Of the existing architectures, most are sequence-to-sequence models; this approach makes it intractable to score (or even enumerate) the entire answer set. R2E perhaps bears the most resemblance to kNN-LM (Khandelwal et al., 2019) which uses retrieval to improve next-token prediction. However, kNN-LM uses retrieval to augment a standard masked language model, while R2E is fully retrieval-based to enable evidence-driven explanations. The Fusion-in-Decoder (FiD) approach (Izacard & Grave, 2021) also bears a notable resemblance to R2E; both merge each piece of evidence with the query independently before jointly processing evidence. While FiD is motivated by efficiency and performance, we are additionally motivated here by explainability.
2.2 Explainability
R2E is inspired by SHAP (SHapley Additive exPlanations) (Lundberg & Lee, 2017), which explains model predictions by approximating Shapley values (Shapley et al., 1953). SHAP evaluates how input features impact model predictions; in part, by hiding subsets of features. Here, we are instead interested in how evidence drives the model prediction. While Data Shapley and related methods (Ghorbani & Zou, 2019; Pandl et al., 2021) have examined the Shapley values of training data, they focus on the impact of training examples on overall model performance rather than inference-time predictions. In fact, evaluating data Shapley values for predictions is typically impossible with fully parametric approaches; after training data has been compressed into the model parameters, there is no straightforward way to hide it for an arbitrary prediction. R2E instead uses a semi-parametric approach where evidence contributes to the prediction at inference time; by turning evidence into features, we can hide evidence and re-evaluate the prediction.
Alternatively, SimplEx (Crabbé et al., 2021) explains predictions by approximating a classified input in terms of a corpus of classified exemplars. SimplEx is general-purpose but indirect: the corpus illuminates black-box model predictions, but does not impact them. In contrast, the corpus directly drives model predictions in R2E, and changing the corpus will change predictions.
2.3 Models for Hypothesis Generation
The use of models in generating or evaluating scientific hypotheses is an emerging area of research. Knowledge graphs are a popular approach for novel hypothesis generation, because their structure enables multi-hop inference between unconnected nodes. Novel hypotheses have been generated by subject-area experts directly querying and inspecting a knowledge graph (Smith et al., 2021).
Sourati & Evans (2023) use knowledge graph patterns for material property prediction and drug re-purposing, additionally leveraging nodes for specific researchers to infer which discoveries are more or less likely to be discovered based on social dynamics. Paliwal et al. (2020) used tensor factorization on a biomedical knowledge graph to predict future research findings and clinical trial outcomes for therapeutic drug targets. Aliper et al. (2023) similarly employed a biomedical knowledge graph to predict clinical trial outcomes; they used a graph transformer network, and ensembled it with a tabular model leveraging clinical trial design features. R2E differs from these approaches by enabling explainability of predictions from the evidence and operating directly on published research without the need to construct a knowledge graph.
In this vein, Tshitoyan et al. (2019) work with a materials science research corpus to identify new material properties. They use cosine similarity on unsupervised word embeddings, specifically word2vec (Mikolov et al., 2013). This approach resembles our parametric masked language model baseline, except that in our case embeddings are derived using a transformer. Tshitoyan et al. suggest that word2vec enables indirect inference similar to that in a knowledge graph; for instance, a material that has never been defined as thermoelectric may nonetheless be mentioned alongside properties associated with thermoelectricity. We observe a similar phenomenon in R2E: for instance, a target never identified directly with a disease may still have been shown to regulate disease-relevant mechanisms (Figure 1) or to be genetically associated with relevant traits (Appendix T). R2E can use these indirect findings as support.
3 Methods
We consider the problem of scoring $N$ potential answers $\mathcal{A}=\{a_{i}\}_{1}^{N}$ to a user query $q$ such that they can be ranked from most to least relevant. We restrict $q$ to be a cloze-style statement as opposed to a natural question (e.g. [MASK] is a promising drug target for the treatment of osteoporosis.), where each answer $a_{i}$ represents a potential named entity for the mask location. Lewis et al. (2019) provides an approach to translate between cloze- and natural-style questions.
Figure 2: R2E architecture schematic. Illustration of R2E inference and explanation. Input: A user-defined cloze-style query, an entity to evaluate, and a corpus of evidence passages corresponding to that entity with entity mentions replaced with [MASK]. Retriever: The query text is encoded with a transformer. All of the entity’s evidence passages are encoded prior to inference, using the same encoder, and stored in a FAISS search index. The $k$ evidence passages with highest cosine similarity to the query are retrieved. Reasoner: The query/evidence pair embeddings are layer-normalised. Each pair is combined at corresponding dimensions into a single embedding using convolutional layers. All combined pair embeddings are passed to a set transformer, followed by a linear layer and sigmoid to obtain the binary probability. Shapley values for each pair (corresponding to each piece of evidence) are computed. Output: To rank over a set of entities, binary probabilities are obtained independently for each entity as described above. Shapley values attribute model predictions back to the evidence passages providing an explanation of the model’s prediction.
3.1 Masked Entity-Linked Corpus
Our approach uses a training corpus of textual passages, $\mathcal{D}$, each containing at least one named entity from the answer set $\mathcal{A}$. An entity linking methodology identifies named entities in $\mathcal{A}$ in the corpus.
For each passage, the span of every occurrence of the named entity is replaced by a [MASK] token. When multiple unique answers occur, we generate a new passage with each masked in turn while the others appear as plain text. Each example is therefore a tuple $(a,d)$ consisting of an entity identifier and a masked text passage in which that entity occurs.
In our specific use case of drug target identification, $\mathcal{A}$ consisted of 19,176 protein-coding gene entities, hereafter referred to collectively as Genes. Surface forms of the gene and corresponding protein (encoded by the gene), were grounded to the same entity under a 1:1 assumption. $\mathcal{D}$ was a large-scale corpus of 160 million sentences from scientific literature. For more details on the corpus and splits used, including temporal splits, see Appendix B; for details on the entity linking methodology, see Appendix C. Note that $|\mathcal{D}|$ could support other use cases (e.g. biomarker identification, drug repurposing, biological mechanism selection) by adjusting the named entity answer set.
3.2 Masked Language Model (MLM)
We first consider a parametric approach to the problem based on the pre-training method in Brayne et al. (2022). We use an encoder-only transformer (Vaswani et al., 2017), specifically a scaled-down version of PubMedBERT (Gu et al., 2021). For query passages $d^{q}\in\mathcal{D}$ containing a masked answer $a_{i}\in\mathcal{A}$, we train the model to predict $p(a_{i}|d^{q})$.
The query embedding is the mean over output embeddings corresponding to [MASK] tokens when the entity is mentioned multiple times. We take the dot product with a learned embedding for each possible answer in the answer set, then apply a bias and softmax to predict $p(a_{i}|d^{q})\,\forall\,a_{i}\in\mathcal{A}$. We train with a cross-entropy loss. Pre-trained weights for domain-specific models are available (e.g. PubMedBERT), but we train from scratch to avoid leakage from pre-training data in our literature evaluations.
This model is both a baseline (MLM) and the basis for the Retriever component of R2E (Section 3.3).
3.3 R2E Retriever
We now consider our semi-parametric approach, R2E, which leverages retrieval from an evidence corpus. R2E combines a Retriever module and a Reasoner module (Figure 2). See Appendix D for additional details of the R2E architecture, training and inference hyperparameters.
The MLM in Section 3.2 has text embeddings that are trained to have high vector similarity with the paired answer embeddings in the answer set. We reasoned that two text embeddings would therefore have high vector similarity if they permit a similar distribution over answers, i.e. if they were semantically similar with respect to this task. This makes the MLM particularly well-suited for identifying corpus passages that are relevant to the user query. We therefore used this MLM for the Retriever module of R2E.
We first used the MLM to embed each of the masked evidence passages in the evidence corpus $\mathcal{D}^{e}$, where $\mathcal{D}^{e}\subset\mathcal{D}$. We partitioned the evidence embeddings according to the masked answer entity they contained, and created distinct FAISS search indices (Johnson et al., 2019) for each.
At inference time, the user’s cloze-style query is encoded with the MLM, yielding $q$. The Retriever selects the $k$ evidence passages $[d^{e}_{1i},...,d^{e}_{ki}]\subset\mathcal{D}^{e}$ with the highest cosine similarity to $q$ from each answer $a_{i}$’s search index. For our experiments, $k=64$. The query and retrieved evidence for each answer are then the inputs to the Reasoner.
3.4 R2E Reasoner
3.4.1 Training Objective
We train the Reasoner with a binary cross entropy loss to differentiate positive examples ($L=1$) from negative examples ($L=0$), i.e. to learn $p(L=1|a_{i},d^{q})$ when taking an entity $a_{i}$ and masked query $d^{q}$ as input. Positive examples were constructed from pairs ($a_{p}$, $d^{q}$) in $\mathcal{D}^{q}\subset\mathcal{D}$, the data split of Reasoner training queries. For each positive example, a corresponding negative example ($a_{n}$, $d^{q}$) was constructed by uniformly sampling $a_{n}\in\mathcal{A}$. For each ($a_{i}$, $d^{q}$) pair, positive or negative, the Reasoner receives the top $k$ evidence passages $[d^{e}_{1i},...,d^{e}_{ki}]$ fetched by the Retriever from the retrieval corpus of $a_{i}$. Under this negative sampling scheme, the objective $p(L=1|a_{i},d^{q})$ is closely related to the MLM multiclass objective $p(a_{i}|d^{q})$ at optimality (Appendix F); however, unlike multinomial regression, sampling negative examples avoids retrieving evidence for every possible alternative answer during training.
3.4.2 Inference
At inference time, we generate $p(L=1|a_{i},q)\,\forall\,a_{i}\in\mathcal{A}$ to score and rank the full answer set for the cloze-style query $q$, using the evidence fetched for $q$. This requires $|\mathcal{A}|$ nearest neighbour searches and forward passes through the Reasoner; however, since retrieval and reasoning for each answer is independent of all other answers, the process can be fully parallelized subject to computational resources.
3.4.3 Architecture
The R2E Reasoner architecture is shown in Figure 2.
[…]
Next, the evidence combiner $g:(\mathbb{R}^{h})^{k}\rightarrow[0,1]$ merges the resulting $k$ query-evidence embeddings and generates $p(L=1|a_{i},q)$. As there is no inherent ordering among the $k$ vectors, we use a set transformer (Lee et al., 2019).
Since the Reasoner sees the answer $a_{i}$ only indirectly via the evidence embeddings, we can also view the score as the probability that the query and evidence embeddings are discussing the same entity.
3.5 R2E Explanations
R2E provides explanations in the form of Shapley values (Shapley et al., 1953; Lundberg & Lee, 2017) attributing the model prediction back to pieces of retrieved evidence, such that they sum up to the overall model score for the query.
Multiple methods exist for rapidly approximating Shapley values on deep learning features (Lundberg & Lee, 2017). Defining each of the $k$ inputs to the evidence combiner as a distinct feature gives a relatively small feature space. As a result, it is tractable to use a simple permutation sampling approach to approximate Shapley values (see Appendix E).
During training, we replaced query-evidence features at random with a learned NULL embedding. In addition to acting as a regularizer (akin to dropout), introducing the NULL embedding during training ensured that the model could handle missing features robustly when estimating Shapley values. For each training example, a dropout rate was sampled in $\text{Uniform}(0,1)$.
3.6 Post-hoc Frequency Bias Correction
Many potential answer sets suffer from class imbalance. In drug target identification, some targets are significantly more well-studied than others. As a result, the learned model $p(a_{i}|q)$ can be strongly correlated with the prior $p(a_{i})$.
While bias can be informative (e.g. reflecting the fact that some targets are involved in more diseases than others) it can also be misleading (e.g. reflecting publishing trends rather than underlying biology). To flexibly control for bias, we therefore use a corrected output based on the frequency of each answer in the training corpus and parameterized by $c\in[0,1]$ (Appendix G). When $c=0$ the scores and rankings are unaltered; when $c=1$, the rankings reflect the pointwise mutual information (PMI) of the query and answer, inspired by the use of PMI in NLP co-occurrence statistics (Church & Hanks, 1990). In the results we report both uncorrected ($c=0$; R2E-uncor) and partially corrected ($c=0.5$; R2E-cor) rankings, with the latter value selected using a validation set (Appendix D). In Shapley value explanations, the bias correction can be represented as an additive feature.
4 Experiments and Results
We evaluate R2E performance on three datasets aligned with drug target identification, each involving prediction over Genes:
  • [itemsep=0.0pt,topsep=0pt,partopsep=0pt]
  • Held-out Biomedical Literature: Predicting masked genes in biomedical literature sentences taken from abstracts published after the publication period of the training data and retrieval corpus.
  • Gene Description Facts: Predicting masked genes in sentences containing human-curated information about the gene, based on gene descriptions provided by UniProt (Consortium, 2022).
  • Clinical Trial Outcomes: Retrospectively predicting success or failure in historical clinical trial data based on evidence published before the trials, using the disease indication and drug target (gene).
For Gene Description Facts and Clinical Trial Outcomes, we also construct Evidence Annotations datasets to evaluate the alignment of R2E explanations with expert reasoning. Specifically, we look at the strength of relationship between R2E Shapley values and GPT-4 (Achiam et al., 2023) binary annotations of whether each piece of explanatory evidence is relevant or irrelevant to the query. We validate GPT-4 annotations against human drug discovery expert annotations.
Evaluation dataset statistics are summarised in Appendix H, and information on accessing datasets in Appendix A.
Table 1: Held-out Biomedical Literature: Ranking metrics on a dataset consisting of one sentence per gene in Genes for each year of 2020, 2021, and 2022. MLM and R2E trained on data published prior to 2020. MCS, R2E-Uncorrected and R2E-Corrected also retrieved data published prior to 2020. R2E-Corrected-Updated retrieved up to the year before the publication year of the query sentence.
Metric Query year Baselines R2E
Freq MCS MLM Uncor Cor Cor-Updated
2020 $<$0.001 0.182 0.181 0.198 0.233 -
MRR 2021 $<$0.001 0.172 0.169 0.187 0.215 0.223
2022 $<$0.001 0.167 0.164 0.178 0.205 0.219
2020 7661 3280 3465 2803 2489 -
MR 2021 7834 3568 3789 3032 2695 2544
2022 7931 3770 4016 3287 2902 2623
2020 $<$0.001 0.268 0.269 0.291 0.333 -
h@10 2021 $<$0.001 0.251 0.252 0.274 0.313 0.324
2022 $<$0.001 0.243 0.243 0.260 0.295 0.312
2020 0.014 0.443 0.438 0.484 0.521 -
h@200 2021 0.014 0.422 0.416 0.456 0.497 0.509
2022 0.013 0.404 0.398 0.435 0.473 0.496
4.1 Metrics
For ranking Genes on Held-out Biomedical Literature and Gene Description Facts, we report mean reciprocal rank (MRR), mean rank (MR), hits@10 (h@10) and hits@200 (h@200). In the case of Gene Description Facts, we used macro ranking metrics to ensure each gene is given equal weight irrespective of frequency. For Clinical Trial Outcomes we report AUROC. For Evidence Annotations, we report AUROC for the R2E Shapley scores of evidence sentences against GPT-4 annotations, and accuracy when validating GPT-4 against human expert annotations.
4.2 Baselines
In addition to MLM, we include two baselines throughout: FREQ and MCS. For FREQ, entities were scored according to their frequency in the training set of $\mathcal{D}$. For MCS (mean cosine similarity), each entity $a_{i}$ was scored by computing $\frac{1}{64}\sum_{j=1}^{64}(d^{e}_{ji}\cdot q)/(\|d^{e}_{ji}\|\|q\|)$ for the query $q$.
For Clinical Trial Outcomes, we also include a genetics baseline, described in detail in Appendix P.
4.3 Predicting Genes in Held-out Biomedical Literature
4.3.1 Dataset Construction
For all experiments in this section, we trained the MLM (R2E Retriever) and R2E Reasoner only on biomedical literature data published prior to 2020. Except where specified, R2E also only retrieved data published prior to 2020 (Figure S3). We then constructed a Held-out Biomedical Literature evaluation dataset from publicly-available paper abstracts. We generated a balanced dataset by sampling one sentence per unique gene in Genes for each of the years 2020, 2021, and 2022. See Appendix B for more detail.
4.3.2 Results
R2E improved on the baselines over all year subsets, both with and without bias correction (Table 1). Bias-corrected R2E improved on uncorrected performance, consistent with the use of a balanced evaluation dataset. We show results on an imbalanced dataset (without stratification by gene in Genes) in Appendix I, demonstrating comparable performance between R2E, MCS and MLM.
To test R2E’s ability to leverage retrieved literature that it was not trained on, we enabled retrieval up to the year preceding the query sentence publication (rather than strictly prior to 2020). This improved performance (Table 1).
We performed two ablation studies: replacing the R2E Retriever with a pre-trained PubMedBERT model (Gu et al., 2021), and replacing the query-evidence encoder of the R2E Reasoner with a Hadamard product $q\circ d^{e}$. R2E outperformed both ablations (Appendix J).
We additionally investigated restricting retrieval to literature data published prior to the query sentence during training, with little consistent impact on performance (Appendix K).
4.4 Predicting Genes in Gene Description Facts
4.4.1 Dataset Construction
We extracted descriptions of protein functions for our Genes entity set from UniProt (Universal Protein Resource) (Consortium, 2022). Each description is a human-written summary of function, and consists of one or more sentences. We used a combination of regular expressions and prompting the Claude 2.0 language model from Anthropic [^1] https://www.anthropic.com/ to extract a set of [MASK]-containing facts from each description. Further details of the source and preprocessing of the dataset, including the Claude prompt and an example description and facts, are found in Appendix L. R2E was trained on, and retrieved from, all years of literature evidence for the Gene Description Facts evaluation.
We also constructed an Evidence Annotations dataset by having GPT-4 annotate (as query relevant or irrelevant) all evidence for 50 randomly sampled Gene Description Facts query-entity pairs (positive examples), and the same 50 queries with randomly sampled alternative entities (negative examples). See Appendix M for the prompting strategy and Appendix N for illustrative annotations. A human drug discovery expert annotated all evidence sentences for a subset of 8 randomly sampled examples (4 positive, 4 negative), following equivalent instructions to the GPT-4 prompt, to validate GPT-4 annotations.
4.4.2 Results
R2E substantially improved on all baselines, both with and without bias correction (Table 2). Bias correction again outperformed the uncorrected model. R2E metrics appear to show greater improvement over baselines than for the Held-out Biomedical Literature dataset. Additionally, there was a strong correlation between evidence Shapley values and GPT-4 relevance annotations, with an AUROC of 0.824. GPT-4 had an accuracy of 0.715 when evaluated against the human-expert annotations.
Table 2: Gene Description Facts: R2E macro ranking metrics on cloze-style queries adapted from human-curated gene descriptions.
Metric Baselines R2E
Freq MCS MLM Uncor Cor
MRR $<$0.001 0.176 0.167 0.202 0.260
MR 8252 1776 2208 937 599
h@10 $<$0.001 0.309 0.296 0.349 0.434
h@200 0.013 0.622 0.590 0.701 0.776
4.5 Predicting Clinical Trial Outcomes
4.5.1 Dataset Construction
We constructed a benchmark of gene-disease pairs (therapeutic hypotheses) with success or failure outcomes in clinical trials as per Nelson et al. 2015 (Appendix O), using the PharmaProjects database (Citeline, ). This benchmark focuses on in vivo efficacy of therapeutic hypotheses as demonstrated by the transition of drugs from Phase II/III clinical trials to regulatory approval.
As genetics data is widely acknowledged to predict successful clinical development (Nelson et al., 2015; Ochoa et al., 2022), we use the most recently published PharmaProjects-aligned genetics dataset (Minikel et al., 2023) (Appendix P) as a competitive baseline. In order to validate our Clinical Trial Outcomes data, we corroborated the published result that the probability of clinical success of therapeutic hypotheses supported by genetics evidence is approximately double the probability without supporting genetics evidence (Relative Success: 1.98; 95% CI (1.76, 2.24)).
We also constructed an Evidence Annotations dataset by having GPT-4 assess the relevance of all evidence for 100 Clinical Trial Outcome therapeutic hypotheses associated with success (50 success, 50 failure; randomly sampled). See Appendix Q for the prompting strategy and Appendix R for illustrative annotations. A human drug discovery expert annotated all evidence sentences for the 8 therapeutic hypotheses for which they had most specialised knowledge (4 success, 4 failure), following equivalent instructions to the GPT-4 prompt, to validate GPT-4 annotations.
To avoid temporal leakage due to the reporting of clinical trial results in the literature, we trained the Reasoner and Retriever with pre-2005 literature (Section 3.1) and filtered out drugs with dated evidence prior to 2005 (Appendix O). We predicted therapeutic hypotheses using the query template “[MASK] is a promising drug target for the treatment of {DISEASE}.”, substituting {DISEASE} with the disease annotated in PharmaProjects (Appendix S).
4.5.2 Multimodality via Templating into Natural Language
We assessed R2E’s ability to reason from genetics data by generating a sentence for every row in the genetics dataset (77,645 total), with the simple template “[MASK] is genetically associated with {MeSH name}.”. The MeSH name, as supplied in Minikel et al. 2023, was programmatically reformatted to better align with naming conventions in the biomedical literature (details in Appendix S). The genetics corpus was given to the R2E Retriever alone and in combination with the year-split biomedical literature.
4.5.3 Results
Table 3: Clinical Trial Outcomes: AUROC for R2E retrieving from literature-alone, genetics-alone, or both; in comparison to baselines.
Model Modality AUROC
Genetic Genetics 0.545
Freq Literature 0.561
MCS Literature 0.623
MLM Literature 0.630
R2E-cor Literature 0.632
R2E-uncor Genetics 0.579
R2E-cor Combined 0.633
R2E-audited-cor Combined 0.638
Primary results are shown in Table 3, with further results across temporal training, bias correction, and method of combining modalities available in Appendix T. Previous analyses restrict clinical trial outcomes to those diseases with genetics data (Minikel et al., 2023); as literature evidence is not restricted to those diseases, we instead show performance against all diseases in the Clinical Trial Outcomes data. Nonetheless, results are qualitatively similar on the restricted set (Table S9).
When reasoning from templated genetics evidence alone, we found that R2E displays improved performance over the genetics baseline despite leveraging the same data. However, R2E had much stronger performance when using the biomedical literature. There was only a marginal improvement from combining templated genetics evidence and the biomedical literature (see additional approaches to combining data sources, with similar performance, in Table S10). There was variance in relative performance across disease areas and modalities (Figure S4).
Shapley values were correlated with GPT-4 relevance annotations, albeit less strongly than in Section 4.4.2, with an AUROC of 0.665. GPT-4 had an accuracy of 82.2% compared to human expert annotations.
4.6 Auditing Explanation Evidence with GPT-4
4.6.1 Dataset Construction
We sought to assess the hypothesis that R2E explanations could enable human- or LLM-in-the-loop feedback to remove false positives. Pooling R2E predictions on the Clinical Trial Outcomes dataset, we used GPT-4 to annotate the relevance of 20,000 query-evidence pairs with the highest Shapley values (computed on pre-sigmoid outputs). We then reran R2E inference on the full dataset, replacing evidence labelled as irrelevant with the NULL embedding. See Appendices Q and U for the GPT-4 prompt and examples.
4.6.2 Results
Removing irrelevant sentences from the coalitions used for R2E predictions resulted in a small but significant increase in AUROC (Table 3, de Long test, $p=0.004$).
In addition, 527/809 high Shapley value genetics sentences were annotated as relevant, and 268/527 of these related to a disease that was neither a substring of, nor contained, the Clinical Trial Outcomes disease.
5 Limitations and Future Work
Retrieving evidence at inference time to make predictions has a cost: generating a score for each answer requires a vector search over the answer’s evidence, followed by a model forward pass. In comparison, predicting from a multiclass model (like the MLM) requires a single forward pass without retrieval. However, retrieval and reasoning are both independent across possible answers, so R2E scales effectively with computational resources.
Retrieval-based inference also has flexibility benefits beyond those explored here. By filtering retrieved evidence on document metadata, users could customize the ranking at inference time; with a scientific literature dataset, this could include filtering supporting evidence to specific timespans, publications, impact factors, paper sections, or keywords.
In Sections 4.4 and 4.5, we applied the model directly to downstream tasks; in the case of clinical trials, we simply adopted a one-size-fits-all query template. Instead, the system could be fine-tuned for the task of interest. Fine-tuning is of particular interest in a human-in-the-loop setting, where a user could identify faulty evidence use (as opposed to a faulty prediction) and tailor the feedback accordingly. Similarly, an LLM could be used to generate labels for model fine-tuning in addition to inference-time auditing.
The evidence templating approach used for genetics is relatively general, and could be applied to other data modalities (e.g. transcriptomics evidence in drug discovery). However, care should be taken with respect to the distribution of the training data (e.g. in drug discovery, templating evidence consistently with how it might be discussed in the literature).
Finally, given the small size of the Retriever and Reasoner architectures (10 and 2 million parameters respectively), we expect improvements with model scaling. Similarly, longer literature passages (e.g. paragraphs rather than sentences) are likely to provide more context and improve performance.
6 Discussion
By retrieving evidence to make predictions, R2E enables explainability, a critical feature in complex, high-risk decision-making. R2E also outperformed the target identification baselines after accounting for class imbalance. Combined with the proposed bias correction technique, this improves the ability to make informed predictions about novel and understudied, but promising targets. Finally, R2E outperformed competing approaches on the difficult task of predicting clinical trial efficacy outcomes, without task-specific fine-tuning. Performance was further improved by auditing R2E’s prediction explanations using GPT-4, an approach made possible by the retrieval-based setup.
When R2E reasoned from genetics data templated into natural language, we found that it significantly outperformed existing approaches (Nelson et al., 2015) leveraging the same data. This could be explained by the language model’s capacity to leverage “soft” semantic associations (e.g. recognizing correlations between diseases beyond those identified by ontological similarity of diseases), as corroborated by the inspection of high-scoring genetics evidence (see Appendix T and Figure S5 for further analysis). This is particularly important given the relatively poor performance of the genetics dataset across all clinical outcomes; while direct genetics evidence is predictive when it exists, it often does not exist or is not applicable. In contrast, the scientific literature has broad coverage across a wide range of indications.
The recent explosion in retrieval-based approaches has challenged the paradigm of using a single model to both memorize factual knowledge and reason from that knowledge. We show here that retrieval can convey not only performance and flexibility advantages, but also significantly improved transparency into how the model reasons from evidence.
Impact Statement
In general, this paper presents work whose goal is to advance the field of Machine Learning. There are many potential widely discussed societal consequences of developments in machine learning, that we do not think warrant specifically highlighting here.
In this paper, we demonstrate the application of R2E to predict and explain protein-coding genes in response to a user query. This application is quite different to either the generality of large language models or the structural biology and chemistry foci of the AI-enabled biological tools most typically associated with any potential dual risk concern. As with other tools that facilitate biomedical research and understanding, the ability to identify and understand particular genes could be applied in a range of use cases. For this paper, we do not believe there to be material risks to highlight, especially noting: (1) We are not releasing proprietary training data, code, or model weights; (2) Explanations provided by R2E are either publicly-available extracts from the scientific literature or structured data templated in natural language, and can be interpreted by expert users in the context of their wider biomedical understanding, but do not significantly lower the barrier to entry for non-experts users; (3) R2E is predicting at the level of drug targets, with multiple complex downstream steps required to translate the identification of a target that may achieve a particular biological effect, into a capability to intervene on that target.
Principles for the ethical deployment of AI at BenevolentAI are outlined on our website.
Acknowledgements
The authors would like to thank Nicola Richmond and Julien Fauqueur for their helpful comments and feedback during drafting, Bradleigh Whitton and James Grey for applying their biological expertise to annotate evidence for validation of GPT-4 Evidence Annotations datasets, Alison McGarvey and Eryk Kropiwnicki for their help and advice on the process for templating genetics data into sentences, Antonios Poulakakis Daktylidis for their assistance with validation of the Clinical Trial Outcomes data, and Hao-Chih Lee for their helpful insights on evaluations of evidence explanations.
Appendix A Accessing Evaluation Datasets
We make the three performance evaluation datasets used in this paper (see Section 4) publicly available, licensed under CC BY-NC-SA 4.0, at: https://github.com/BenevolentAI/r2e-evaluation-data. Specific licensing information for the datasets is as follows:
  • Clinical Trials Outcomes © 2024 by BenevolentAI is licensed under CC BY-NC-SA 4.0. We have permission from Citeline PharmaProjects to publicly release the subset of their data that is used here.
  • Gene Description Facts © 2024 by BenevolentAI is licensed under CC BY-NC-SA 4.0. It is adapted from ”Universal Protein Resource (UniProt)” by Uniprot Consortium, used under CC BY 4.0.
  • Held-out Biomedical Literature validation and test dataset sentences are courtesy of the National Library of Medicine.
Appendix B Masked Entity-Linked Corpus, Dataset Splits & Sizes
The large-scale corpus of scientific documents consisted of open access PubMed abstracts and PMC full texts as well as paid access Springer, Wiley and Elsevier full texts. We performed entity linking using a proprietary method (Appendix C), however any entity linking approach may be used (e.g. dictionary matching). Individual sentences were used as passages.
We filtered to sentences in the corpus that contained both: i) one or more protein-coding genes (entity set referred to as Genes), and ii) one or more non-gene grounded biomedical entities (e.g. diseases, biological pathways etc.), to select for an informative corpus. This process yielded 160 million sentences.
We created three distinct corpus splits $\mathcal{S}_{1}$, $\mathcal{S}_{2}$, and $\mathcal{S}_{3}$ to handle training, validation and testing across the both MLM/Retriever and Reasoner (Figure S3). These splits were generated at the level of entire documents to reduce the occurrence of highly similar sentences between splits.
For the MLM/Retriever, datasets were composed as follows:
  • $\mathcal{D}^{\mathrm{MLM}}_{train}=\mathcal{S}_{1}$
  • $\mathcal{D}^{\mathrm{MLM}}_{val}=\mathcal{S}_{2}$
For R2E, we use $\mathcal{D}^{e}$ to refer to an evidence/retrieval corpus and $\mathcal{D}^{q}$ to refer to a query corpus. The datasets were composed as follows:
  • $\mathcal{D}^{e}_{train}=\mathcal{S}_{1}$
  • $\mathcal{D}^{q}_{train}=\mathcal{S}_{2}$
  • $\mathcal{D}^{e}_{val}=\mathcal{S}_{1}\cup\mathcal{S}_{2}$
  • $\mathcal{D}^{q}_{val}\subset\mathcal{S}_{3}$
  • $\mathcal{D}^{e}_{test}=\mathcal{S}_{1}\cup\mathcal{S}_{2}$
  • $\mathcal{D}^{q}_{test}\subset\mathcal{S}_{3}:\mathcal{D}^{q}_{test}\cap\mathcal{D}^{q}_{val}=\emptyset$
$\mathcal{S}_{3}$ was partitioned from $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, either randomly (with 1.5 million sentences allocated to $\mathcal{S}_{3}$) or using a year split (all documents from a specified year onwards allocated to $\mathcal{S}_{3}$ and all documents from prior years allocated to $\mathcal{S}_{1}$ or $\mathcal{S}_{2}$). $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ were always created by a random partition with 1.5 million sentences allocated to $\mathcal{S}_{2}$ and the remainder in $\mathcal{S}_{1}$.
The number of sentences in $\mathcal{S}_{1}$ for each of the year splits used in our reported experiments, corresponding to particular evaluations, were:
  • 2005 year split (${\mathcal{D}}^{q}_{eval}=$ Clinical Trial Outcomes): 16.2 million
  • 2020 year split (${\mathcal{D}}^{q}_{eval}={\mathcal{D}}^{q}_{test}=$ Held-out Biomedical Literature): 112 million
  • Full data, no year split (${\mathcal{D}}^{q}_{eval}=$ Gene Description Facts): 157 million
The above splitting procedure is illustrated in Figure S3 for the case of a 2020 year split model. In the 2020 year split case, note the disjoint subsets sampled from $\mathcal{S}_{3}$ and used to create overall validation ($\mathcal{D}^{q}_{val}$) and test ($\mathcal{D}^{q}_{test}$) queries, for ranking metric evaluations over all genes in Genes; namely the:
  • Held-out Biomedical Literature validation dataset: 1 sentence per gene, sampled from publicly-available abstract section sentences from 2020 onwards. Used for hyperparameter selection and ablation experiments described in Appendices D & J respectively.
  • Held-out Biomedical Literature test dataset: 1 sentence per gene per year for 2020 onwards, sampled from publicly-available abstract section sentences. Used for evaluations described in Section 4.3.1, including evaluation of the MLM and other baselines.
Figure S3: Masked entity-linked corpus data splits with a 2020 split year. Here we illustrate how the masked entity-linked corpus was partitioned to enable MLM/Reasoner and Retriever training, validation, and testing. Specifically the example of a 2020 year split dataset is shown, with overall validation and testing of R2E (and all baselines) using queries from 2020 onwards ($\mathcal{D}^{q}_{val}$ and $\mathcal{D}^{q}_{test}$, disjoint subsets sampled from Split 3). The overall validation and test query sets here correspond to the Held-out Biomedical Literature validation and test datasets. The partitioning setup shown here for the 2020 year split, is illustrative of other data split setups used in our experiments - the key differences being that the Clinical Trial Outcomes evaluation used a 2005 year split dataset, and Gene Description Facts evaluation used a random split rather than a year split dataset.
Appendix C Entity Linking
We used a proprietary entity linking methodology based on dictionaries of entities and synonyms, as well as an abbreviation detection algorithm and a model that resolves type ambiguities based on the context of each mention. The dictionaries were created from several sources.
  1. 1.
    External ontologies.
  2. 2.
    Human annotations of synonyms discovered by machine learning methods.
  3. 3.
    Automatic synonym generation to cover e.g. variation in punctuation, Greek letters and plurals of terms.
For the protein-coding gene target entities, referred to as Genes and used throughout the paper, we ground both gene and protein forms to the same entity, under the assumption of a 1:1 relationship between a gene and the protein it encodes.
Appendix D R2E Hyperparameters
The R2E model was implemented using PyTorch deep learning library (Paszke et al., 2019).
All sentences were tokenized, and then truncated and padded to a length of 128, using the same vocabulary as PubMedBERT (Gu et al., 2021). Pre-processing of training examples for both Retriever and Reasoner training was done using Apache Spark (Zaharia et al., 2016). Both the Retriever and Reasoner were trained independently, for 10 epochs on a single Tesla V100 GPU, with a total training time of approximately 1 week.
The final R2E Retriever architecture, as well as the MLM baseline, consisted of a scaled down version of PubMedBERT (Gu et al., 2021) trained from scratch on the task described in 3.2, with 4 hidden layers, 4 attention heads, an intermediate size of 512, a hidden size of 256, and total size of 10 million parameters. Final Retriever/MLM training used a batch size of 512, a categorical cross-entropy loss, and an AdamW optimizer (Loshchilov & Hutter, 2019) with a learning rate of 0.0001 and no weight decay.
The R2E architecture is summarised in Figure 2. The final query-evidence encoder component of the R2E Reasoner architecture consisted of a layer normalisation across all concatenated query/evidence pairs, then two conv1d layers each with kernel size of 1 (first layer: 2 input channels, 8 output channels; second layer: 8 input channels, 1 output channel) across each query/evidence pair individually. The final evidence combiner component of the R2E Reasoner architecture consisted of a set transformer (Lee et al., 2019) over all query-evidence embeddings returning a single embedding, followed by a linear layer and sigmoid to output a binary probability. The set transformer had 4 heads, 2 induced set attention blocks with 32 inducing points for the encoder, and a pooling by multihead attention followed by two set attention blocks in the decoder. The Reasoner had a total size of 2 million parameters. After freezing the Retriever weights, the final Reasoner training used a batch size of 2048, binary cross-entropy loss, and AdamW optimizer with a learning rate of 0.0001 and weight decay of 0.001. For both training and inference, 64 evidence sentences were retrieved for a given query. A post-hoc frequency bias correction factor of 0.5 was used for the R2E-Corrected variant (Section 3.6 and Appendix G for details of post-hoc correction).
The post-hoc frequency bias correction factor selection and architectural comparison ablations (Appendix J) were based on MRR for a 2020 year split model, on a Held-out Biomedical Literature validation set containing one biomedical literature cloze-style query sentence per gene in Genes from publicly-available abstract sections (see Appendix B). The resulting 15477 validation set queries were therefore sentences published from 2020 onwards, and retrieval corpus sentences published prior to 2020. The learning rate was chosen to reduce training time while maintaining training stability, and the batch size selected to optimise GPU utilisation. We did not evaluate variations of model scale and leave this to future work.
Appendix E Approximating Evidence Shapley Values
We used a simple Monte Carlo method to approximate Shapley values, combined with antithetical sampling for variance reduction (Mitchell et al., 2022). The Shapley value was approximated as
$\phi_{i}\approx\frac{1}{2M}\sum_{j=1}^{M}\left(\left[g(S_{j}\cup\{i\})-g(S_{j})\right]+\left[g(\bar{S}_{j}\cup\{i\})-g(\bar{S}_{j})\right]\right)$ (1)
where $\phi_{i}$ is the approximate Shapley value of feature $i$ (an encoded query/evidence pair), $M$ is the chosen number of sampled permutations, $S_{j}$ is the set of features preceding $i$ in the j-$th$ permutation sample, $g(S_{j})$ is the Reasoner output when only the features $S_{j}$ are unmasked, $g(S_{j}\cup\{i\})$ is the Reasoner output when feature $i$ is unmasked in addition to $S_{j}$, and $\bar{S}_{j}$ corresponds to the set of features preceding $i$ in the reverse of the j-$th$ permutation sample (equivalently, the set of features following $i$ in the j-$th$ permutation sample). The sum of the Shapley values over features plus the score when all features are $\mathrm{NULL}$ equates to the final score. Depending on the purpose, we use either the post-sigmoid output or the logit score for $g$. We use $M=100$ whenever Shapley values are computed as part of this paper.
Algorithm 1 Generate permutation-approximated Shapley attributions for a single query.
  Input: Number of permutations $M$, query-evidence embeddings $E=\{d^{qe}_{1},\ldots,d^{qe}_{k}\}$, missing evidence embedding $NULL$, model forward function $g(\cdot)$
  Output: Shapley value of query-evidence embeddings: $\phi_{1},\ldots,\phi_{k}$
  Initialize $\phi_{i}=0$ for $i=1,\ldots,k$
  $2M$ antithetical sample of permutations $p^{j}$ for $j\in 1,\ldots,2M$ of the feature indices $1,\ldots,k$
  where $p^{M+i}=\mathrm{ReverseOrder}(p^{i})$
  $\tilde{E}_{0}=\{\mathrm{NULL},\ldots,\mathrm{NULL}\}$, with $|\tilde{E}_{0}|=k$
  $s_{0}\leftarrow g(\tilde{E}_{0})$
  for all $j\in\{1,\ldots,2M\}$ do
     for all $i\in\{1,\ldots,k\}$ do
        $\tilde{E}^{j}_{i}\leftarrow\{d^{qe}_{p^{j}[1]},\ldots,d^{qe}_{p^{j}[i]},\mathrm{NULL},\ldots,\mathrm{NULL}\}$, with $|\tilde{E}^{j}_{i}|=k$
        $s_{i}\leftarrow g(\tilde{E}^{j}_{i})$
        $\phi_{p^{j}[i]}\leftarrow\frac{j-1}{j}\phi_{p^{j}[i]}+\frac{1}{j}(s_{i}-s_{i-1})$Cumulative average of marginals for feature $p^{j}[i]$ across permutations}
     end for
  end for
Appendix F Relationship between Multinomial and Binary Objectives
R2E is trained to predict the probability that a given query-entity pair is “true”, i.e. that it came from a real occurrence in the literature and was not randomly generated. Given the labels $L\in\{0,1\}$, the query (masked sentence) variable $Q$, the named entity answer variable $A$, the Reasoner parameters $\theta$ and the fixed Retriever parameters $\psi$, the model is trained to predict
$\frac{1}{1+\exp(-z(a_{i},q_{i}))}\approx P(L=1|Q=q_{i},A=a_{i};\theta,\psi)$ (2)
where $z(a_{i},q_{i})$ is the logit output of the network in response to a specific example $i$, i.e.
$z(a_{i},q_{i})\approx$ $\log(P(L=1|Q=q_{i},A=a_{i};\theta,\psi))$ (3)
-- $\log(P(L=0|Q=q_{i},A=a_{i};\theta,\psi)).$
Here, when $L=0$, the example $i$ corresponds to a negative example where $Q$ and $A$ have been chosen independently. Consider the case where the specific parameters $\theta$ and $\psi$ have been learned such that the equality in Eq. 3 holds exactly; we are interested in the output in this case. We therefore assume the optimal output $z^{*}(a_{i},q_{i})$ and exclude the parameters.
The equation can be re-written using Bayes’ Theorem,
$z^{*}(a_{i},q_{i})=$ $\log(P(Q=q_{i},A=a_{i}|L=1))+\log(p(L=1))-\log(P(Q=q_{i},A=a_{i}))$ (4)
-- $\log(P(Q=q_{i},A=a_{i}|L=0))-\log(p(L=0))+\log(P(Q=q_{i},A=a_{i})).$
In our training setup, positive and negative examples are sampled equally often, i.e.
$\log(p(L=1))=\log(p(L=0)).$ (5)
As a result, Eq. 4 simplifies to
$z^{*}(a_{i},q_{i})=$ $\log(P(Q=q_{i},A=a_{i}|L=1))-\log(P(Q=q_{i},A=a_{i}|L=0))$ (6)
Using the product rule
$z^{*}(a_{i},q_{i})=$ $\log(P(A=a_{i}|Q=q_{i},L=1))+\log(P(Q=q_{i}|L=1))$
-- $\log(P(A=a_{i}|Q=q_{i},L=0))-\log(P(Q=q_{i}|L=0)).$
The distribution over queries is also equal for positive and negative labels, as each query sentence is chosen for each condition once per epoch, simplifying to
$z^{*}(a_{i},q_{i})=$ $\log(P(A=a_{i}|Q=q_{i},L=1))-\log(P(A=a_{i}|Q=q_{i},L=0)).$ (8)
The distribution over named entity answers is independent of the query when conditioned on $L=0$, because negative samples are chosen by randomly pairing queries and entities. So the second term here corresponds to our negative sampling distribution. Therefore, the output at optimality corresponds to
$z^{*}(a_{i},q_{i})$ $=\log(P(A=a_{i}|Q=q_{i},L=1))-\log(P(A=a_{i}|L=0))$ (9)
$=\log(P(A=a_{i}|Q=q_{i},L=1))+\log(|\mathcal{A}|)$
since the probability of choosing a given answer $a_{i}$ as a negative sample during training is $\frac{1}{|\mathcal{A}|}$. Comparing to the optimal logit output of the MLM model, we see a close relationship:
$z^{*,\mathrm{MLM}}(a_{i},q_{i})=$ $\log(P(A=a_{i}|Q=q_{i},L=1))+\log(Z)$ (10)
where $Z$ is the partition function (the MLM includes $L=1$ implicitly as all examples are positive). The optimal logit outputs for the models therefore scale up to their respective normalization factors.
Appendix G Post-hoc Frequency Bias Correction as Trading off Log Probability and Mutual Information
From Equation 9 in Appendix F, we find that the optimal model output logit scales with $\log(P(A|Q,L=1))$, i.e. the probability of the answer given the query assuming a real example ($L=1$). This score will be highly correlated with the prior distribution over the answer set, particularly for an imbalanced dataset (like the mentions of Genes in the scientific literature corpus used in the paper).
One approach to counteract the literature bias, if desired, is to instead consider the pointwise mutual information between a given answer and a given query:
$\mathrm{PMI}(A=a;Q=q)=\log{\frac{P(A=a|Q=q)}{P(A=a)}}.$ (11)
PMI is widely used in the NLP community to measure associations between keywords in a corpus, based on their marginal occurrence counts and joint co-occurrence counts (Jurafsky & Martin, 2019). Similarly, we find that it offers a straightforward means of correcting for class imbalance after training the model.
For a model that predicts a multiclass output (like the MLM), we can directly adapt the output. Specifically, after normalizing the outputs to remove $\log(Z)$, where Z is the partition function,
$z^{\mathrm{MLM}}_{c}(a_{i},q_{i})$ $=z^{\mathrm{MLM}}(a_{i},q_{i})-c\cdot\log P(A=a_{i}|L=1)$ (12)
$\approx\log{P(A=a_{i}|Q=q_{i},L=1)}-c\cdot\log P(A=a_{i}|L=1)$
where $P(A=a_{i}|L=1)$ is estimated by the proportion of passages in the corpus where $a_{i}$ is the correct answer. When $c=0.0$, the two approaches are equivalent; while when $c=1.0$, the output approximates the PMI score in Equation 11. Stronger corrections penalize common answers, and the score is only positive if the model’s estimated answer probability for the given query is higher than the frequency-based prior.
In R2E, we instead note that the optimal logit score in Equation 9 already reflects PMI if the negative sampling probability $P(A=a_{i}|L_{i}=0)$ was chosen to reflect the prior distribution over answers in the dataset, $P(A=a_{i}|L=1)$. We therefore consider a negative distribution $P_{c}(A=a_{i}|L=0)$ that trades off between a uniform distribution $\frac{1}{|\mathcal{M}|}$ and one based on the answer prior in the training corpus:
$P_{c}(A=a_{i}|L=0)=\frac{C(a_{i})^{c}}{\sum_{i=1}^{|\mathcal{A}|}C(a_{i})^{c}}$ (13)
where $C(a_{i})$ is the count of occurrences of answer $a_{i}$ as a masked entity in the training corpus. When $c=1$, this corresponds to the background distribution of $a_{i}$ in the training corpus $P(A=a_{i}|L=1)$; when $c=0$, it corresponds to the uniform distribution $\frac{1}{|\mathcal{A}|}$.
One possible approach to bias correction is to set a fixed $c$ during training and use the resulting negative sampling distribution in Equation 13. However, this approach grants less flexibility in terms of the desired bias correction at inference time. We therefore continue to use the fixed uniform distribution $\frac{1}{|\mathcal{A}|}$ and instead introduce a correction factor
$f_{c}=\log{\frac{1}{|\mathcal{A}|}}-\log{P_{c}(A=a_{i}|L=0)}.$ (14)
Applying this correction to the logit output of R2E after training (Equation 9) yields
$z(a_{i},q_{i})+f_{c}\approx\log(P(A=a_{i}|Q=q_{i},L=1))-\log{P_{c}(A=a_{i}|L=0)}$ (15)
which reflects a log probability estimate when $c=0$ and a pointwise mutual information estimate when $c=1$. We found that the best performance in terms of MRR on the Held-out Biomedical Literature validation dataset (see Appendix B), was achieved with a partial correction of $c=0.5$. We refer to this as R2E-cor, and refer to the case with $c=0.0$ as R2E-uncor.
The bias correction can be straightforwardly identified as an additional additive feature during Shapley value estimation to communicate its impact to the user. For under-represented answers, it can be seen as compensating for “missing” evidence, e.g. due to the lack of research on a particular target.
Appendix H Evaluation Dataset Statistics
The total sizes of all test/evaluation datasets are shown in Table S4.
Table S4: Evaluation dataset statistics
Dataset
Subset Count
Held-out Biomedical Literature 2020 14429
2021 14859
2022 15074
Gene Description Facts
60839
GDF Evidence Annotations (human expert) Query-gene pairs 8
Positives:Negatives 4:4
Evidence 512
GDF Evidence Annotations (GPT-4) Query-gene pairs 100
Positives:Negatives 50:50
Evidence 6400
Clinical Trial Outcomes (2005 onwards) Successes 1449
Fails 4222
CTO Evidence Annotations (human expert) Query-target pairs 8
Successes:Fails 4:4
Evidence 512
CTO Evidence Annotations (GPT-4) Query-target pairs 100
Successes:Fails 50:50
Evidence 6400
Appendix I Comparison of Models on a Non-Stratified Held-out Biomedical Literature Dataset
For completeness, we evaluated R2E on a dataset of 20,000 randomly-sampled publicly-available abstract sentences published from 2020 onwards, obtaining an imbalanced dataset without stratification by gene in Genes. As expected, the frequency-based baseline performs significantly better here relative to the stratified dataset in Table 1. Ranking metrics show similar performance for R2E, MCS and MLM (Table S5). In comparison, on the stratified dataset results in Table 1, R2E markedly outperforms baselines. This suggests that R2E obtains superior performance on the Held-out Biomedical Literature specifically after accounting for class imbalance (improved further by the bias correction procedure) and similar performance when class imbalance is not taken into account.
Table S5: Non-stratified Held-out Biomedical literature: R2E ranking metrics on a random subsplit (not stratified by gene in Genes) of query sentences published from 2020 onwards (20,000 queries), for an R2E model trained and retrieving from data prior to 2020.
Metric Baselines R2E
Freq MCS MLM Uncor Cor
MRR 0.026 0.405 0.399 0.403 0.350
MR 2321 1114 1305 1140 1456
h@10 0.070 0.520 0.519 0.523 0.500
h@200 0.304 0.691 0.686 0.699 0.686
Appendix J Architecture Ablation Experiments
A Held-out Biomedical Literature validation set was used for ablations experiments, consisting of one sentence per gene in Genes sampled from publicly-available abstract sentences published from 2020 onwards (as described in Appendices B & D), for an R2E model trained and retrieving from data prior to 2020. Ablation comparisons were made without applying any post-hoc frequency bias correction. Results are shown in Table S6.
We assessed R2E performance with an off-the-shelf biomedical transformer for the Retriever, in place of our task specific Retriever (Sections 3.2 & 3.3). We ablated the R2E Retriever by substituting in a PubMedBERT model initialised with its published weights (Gu et al., 2021). We created sentence embeddings by taking the mean over outputs corresponding to [MASK] tokens. This Retriever had a larger hidden size with 768 dimensional query and evidence embeddings. The R2E Reasoner was therefore linearly scaled to match this hidden size. We found that the default R2E had better ranking metrics than R2E with PubMedBERT substituted in for the Retriever.
We also assessed a simplification of the R2E Reasoner (Section 3.4.3) by substituting the convolutional query-evidence encoder with a parameter-free Hadamard product between the query embedding and each evidence embedding. The Hadamard product was chosen in order to incorporate an inductive bias towards the cosine similarity. Compared to R2E with the convolutional layers substituted for a Hadamard product, default R2E improved across all ranking metrics.
Table S6: Architecture ablations: Ablated versions of R2E-Uncorrected on a validation dataset consisting of one sentence per gene in Genes sampled from sentences published from 2020 onwards, while training on and retrieving from data prior to 2020. Hadamard: substituting the convolution layers of the Reasoner with a Hadamard product. PubMedBERT: substituting the Retriever for the PubMedBERT model.
Metric R2E-Uncor Hadamard PubMedBERT
MRR 0.181 0.166 0.134
MR 3210 3260 3606
h@10 0.262 0.253 0.207
h@200 0.443 0.441 0.389
Appendix K Temporal Training
We tried a temporal training setup as an attempt to improve R2E’s ability to infer on future (2020 onwards) literature in the Held-out Biomedical Literature dataset from prior (pre-2020) retrieved literature. Specifically, we restricted retrieval during training to biomedical literature evidence sentences published prior to the query training sentence. This resulted in only small shifts in ranking metrics (Table S7).
Table S7: Temporal training: Held-out Biomedical Literature R2E-Corrected ranking metrics when restricting retrieval to biomedical literature published prior to the query sentence, during inference alone verses during training and inference.
Metric Query year R2E-Cor with pre-query retrieval
Inference-only Train & Inference
2020 0.233 0.220
MRR 2021 0.223 0.212
2022 0.219 0.207
2020 2489 2381
MR 2021 2544 2462
2022 2623 2549
2020 0.333 0.325
h@10 2021 0.324 0.315
2022 0.312 0.302
2020 0.521 0.525
h@200 2021 0.509 0.508
2022 0.496 0.493
Appendix L Further Details on Creation of Gene Description Facts Dataset
We downloaded UniProt FTP server data version 2023_01 and extracted descriptions of protein functions for our set of protein-coding gene entities (Genes) from UniProt (Universal Protein Resource), used under CC BY (4.0), (Consortium, 2022) (by pulling “text” from UniProt entities with type “function” in the “comment” field). Each entry is a human-written description of function, and consists of one or more sentences.
After dropping all descriptions containing fewer than four words, we converted each description into a set of single-sentence facts as follows:
  1. 1.
    Descriptions were split into individual sentences and PubMed IDs removed, using regular expression operations.
  2. 2.
    Each sentence was converted into a fact containing a “[MASK]” referring to the gene and “[MASK]” in place of all gene mentions, using one-shot prompted Claude 2.0 language model from Anthropic (prompt template below). Sentences which Claude determined did not contain a suitable fact, were dropped.
  3. 3.
    “[MASK]”-containing facts were extracted from the Claude completion, and facts without any “[MASK]” mention were dropped.
For example, the description for the protein corresponding to gene ELF2 is:
“Isoform 1 transcriptionally activates the LYN and BLK promoters and acts synergistically with RUNX1 to transactivate the BLK promoter. Isoform 2 may function in repression of RUNX1-mediated transactivation.”
From this description, the following facts were extracted for the evaluation dataset:
  • [MASK] isoform 1 transcriptionally activates the LYN and BLK promoters and acts synergistically with RUNX1 to transactivate the BLK promoter.
  • [MASK] isoform 2 may function in repression of RUNX1-mediated transactivation.
The following one-shot prompt template was used to convert sentences from pulled UniProt gene descriptions into [MASK]-containing facts. The gene GENE_NAME and UNIPROT_DESCRIPTION_SENTENCES were substituted into the template for each sentence-gene pair in the dataset, prior to querying Claude 2.0 via Anthropic’s API.
    {HUMAN_PROMPT}
    # THE TASK:
    You are an expert biologist. You will be given a set of sentences from a
    DESCRIPTION of a GENE from UniProt.
    Your instructions are to go one-by-one through each sentence in the
    DESCRIPTION, and:
    1. If the sentence states a fact about the specified GENE convert the
    sentence into a FACT according to the FACT formatting shown in the <example>
    below. 2. If, and only if, the sentence does not state any information
    about the GENE, you may skip the sentence and indicate this with
    "sentence[nb] SKIPPED" as shown in the <example> below.
    # FORMATTING:
    Here’s an example input and output contained in the <example> XML tags,
    to illustrate the format in which FACTs should be stated, including how to
    indicate that a sentence has been skipped.
    <example>
    Input:
    GENE: PGP
    DESCRIPTION sentences:
    <sentence1>Glycerol-3-phosphate phosphatase hydrolyzing glycerol-3-phosphate
    into glycerol.</sentence1>
    <sentence2>Thereby, regulates the cellular levels of glycerol-3-phosphate a
[…]
    <sentence3>Was also shown to have a 2-phosphoglycolate phosphatase activity
    and a tyrosine-protein phosphatase activity.</sentence3>
    <sentence4>However, their physiological relevance is unclear
    (PubMed:26755581).</sentence4>
    <sentence5>In vitro, has also a phosphatase activity toward ADP, ATP, GDP
    and GTP (By similarity).</sentence5>
    <sentence6>Further work is needed to understand this.</sentence6>
    <sentence7>(Microbial infection) Involved in replication of Rubella virus.
    </sentence7>
    Output:
    Here are complete set of [MASK]-containing FACTs for each sentence about PGP:
    <sentence1_fact>[MASK] is a glycerol-3-phosphate phosphatase that hydrolyzes
    glycerol-3-phosphate into glycerol.</sentence1_fact>
    <sentence2_fact>[MASK] regulates cellular levels of glycerol-3-phosphate, a
    metabolic intermediate of glucose, lipid and energy metabolism.
    </sentence2_fact>
    <sentence3_fact>[MASK] has 2-phosphoglycolate phosphatase activity and
    tyrosine-protein phosphatase activity.</sentence3_fact>
    <sentence4_fact>sentence4 SKIPPED</sentence4_fact>
    <sentence5_fact>In vitro, [MASK] has phosphatase activity toward ADP, ATP,
    GDP and GTP.</sentence5_fact>
    <sentence6_fact>sentence6 SKIPPED</sentence6_fact>
    <sentence7_fact>[MASK] is involved in replication of Rubella virus.
    </sentence7_fact>
    </example>
    # FACT REQUIREMENTS
    You must note the following requirements, when constructing each FACT:
    1. Each and every FACT must include one or more [MASK] tokens representing
    the GENE.
    2. All references to or synonyms of the GENE anywhere in a FACT, must also
    be replaced by [MASK].
    3. Only include information explicitly stated in the DESCRIPTION sentence
    when extracting a FACT - do not elaborate with any additional information
    from elsewhere.
    4. You must go through every sentence.
    5. You can only skip a sentence if it contains no information about the
    GENE, and you must indicate this by stating the sentence was SKIPPED in
    the corresponding sentence FACT XML tags.
    # THE FINAL GENE AND DESCRIPTION SENTENCES
    Now, paying attention to all the above instructions and example, please go
    one-by-one through each sentence in the following DESCRIPTION and extract
    each FACT for the stated GENE:
    Input:
    GENE: {GENE_NAME}
[…]
    {AI_PROMPT}
    Output:
    Here are complete set of [MASK]-containing FACT(s) for each sentence about
    {GENE_NAME}:
    <sentence1_fact>
Appendix M Further Details on Creation of Explanation Annotations for Gene Description Facts Dataset
We constructed Evidence Annotations for the Gene Description Facts dataset by having GPT-4 annotate query relevance for all evidence across 50 randomly sampled query-entity pairs (positive examples) and the same 50 queries but with a randomly sampled alternative entity (negative examples), resulting in 6400 query-evidence pairs with a binary annotation.
We used GPT-4 to obtain relevant/irrelevant annotations for this task by using the following prompt, substituting in GENE_DESCRIPTION_FACT and EVIDENCE_SENTENCE:
    You are a scientific expert working on target identification in drug
    discovery.
    Your task is to use your expertise to evaluate whether a piece of evidence
    (referred to as EVIDENCE) about a masked target from an academic paper (in
    the form of a sentence), provides relevant support to a specified biological
    fact about that masked target (referred to as FACT). Please explain your
    reasoning first before giving your answer. Provide your final answer by
    stating either <answer>RELEVANT</answer> or <answer>IRRELEVANT</answer>
    to indicate whether the evidence sentence provides relevant support for
    the FACT or not. Please also indicate your confidence in your answer by
    writing one of <confidence>HIGH</confidence> or <confidence>LOW</confidence>.
    Here are some examples demonstrating proper formatting and reasoning in a
    response:
    <example>
    H: Your FACT of interest is:
    <fact>
    [MASK] also inhibits Wnt signaling by binding to the CTNNB1 protein,
    preventing interaction of CTNNB1 with TCF7L2/TCF4.
    </fact>
    Here is the EVIDENCE sentence to assess:
    <evidence>
    In the HTB-72 melanoma cell line, [MASK] inhibited melanoma cell growth
    and additionally, [MASK] also induced apoptosis.
    </evidence>
    A:The evidence sentence makes no indication that [MASK] is involved in Wnt
    signaling or anything else mentioned in the given biological fact. Therefore,
    it is <answer>IRRELEVANT</answer> with <confidence>HIGH</confidence>.
    </example>
    <example>
    H: Your FACT of interest is:
    <fact>
    [MASK] is a probable serine protease that plays a role in hearing.
    </fact>
    Here is the EVIDENCE sentence to assess:
    <evidence>
    type hearing loss, as occurred in subject SB114-206, caught our attention
    (Figure 1a) because alterations in [MASK], a serine protease, are also known
    to cause progressive ski slopetype hearing loss.
    </evidence>
    A:The evidence sentence states that [MASK] is a serine protease that can affect
    hearing so it directly supports our biological fact being true of [MASK].
    Therefore, it is <answer>RELEVANT</answer> with <confidence>HIGH</confidence>.
    </example>
    Now here is your real task.
    Your FACT of interest is:
    <fact>
    {GENE_DESCRIPTION_FACT}
    </fact>
    Here is the EVIDENCE sentence to assess:
    <evidence>
    {EVIDENCE_SENTENCE}
    </evidence>
Appendix N Examples of Evidence Annotations for Gene Description Facts Dataset
Here we show an example of each of true positive, true negative, false positive and false negative R2E evidence Shapley values versus GPT-4 relevant/irrelevant Evidence Annotations, for Gene Description Facts. These examples also illustrate the level of biological complexity in the Gene Description Facts task.
  1. 1.
    True positive: R2E assigns a large, positive Shapley value to a piece of evidence that GPT-4 labels as relevant. R2E and GPT-4 recognise this as a relevant piece of evidence.
        Query: [MASK] regulates actin dynamics in motor neurons by inhibiting
        the GTP-binding activity of ARF6, leading to ARF6 inactivation.
        Target: C9orf72
        Evidence: The observation that the GTP-bound forms of Rac1 and Arf6 are
        increased when [MASK] is repressed in primary mouse motor neurons or
        in patient iPSC-derived motor neurons strongly suggests that [MASK] is
        essential for the GTP exchange activity of Arf6 in particular.
        (PMID: 27723745)
        R2E-generated Shapley value: 0.0391
        GPT4 completion: A: The evidence sentence directly supports the biological
        fact. It states that when [MASK] is repressed, the GTP-bound forms of ARF6
        are increased, suggesting that [MASK] is essential for the GTP exchange
        activity of ARF6. This aligns with the fact that [MASK] regulates actin
        dynamics in motor neurons by inhibiting the GTP-binding activity of ARF6,
        leading to ARF6 inactivation. Therefore, it is <answer>RELEVANT</answer>
        with <confidence>HIGH</confidence>.
  2. 2.
    True negative: R2E assigns a negative Shapley value to a piece of evidence that GPT-4 labels as irrelevant. R2E and GPT-4 recognise this as an irrelevant piece of evidence.
        Query: [MASK] binds the 3’-UTR of TRPV1 mRNA and represses TRPV1
        translation, which is required to maintain normal thermoception.
        Target: DCTN6
        Evidence: Gene co-expression analyses for SLC2A4 were conducted with
        West Breast [ 28 ], and the results showed that SLC2A4 is co-expressed with
        FADD, BLOC151, RHOB, [MASK], CELF2, SNTB2, NPPB, TIE1, FGFR1, IDH1, ECH1 ,
        etc. (Fig. 5 D). (PMID: 34525987)
        R2E-generated Shapley value: -0.0074
        GPT4 completion: A: The evidence sentence only mentions that [MASK] is
        co-expressed with SLC2A4 and other genes, but it does not provide any
        information about [MASK] binding to the 3’-UTR of TRPV1 mRNA or
        repressing TRPV1 translation. Therefore, it is <answer>IRRELEVANT</answer>
        with <confidence>HIGH</confidence>.
  3. 3.
    False positive: R2E assigns a large, positive Shapley value to a piece of evidence that GPT-4 labels as irrelevant. R2E seems to consider the mention of magnesium concentration in relation to [MASK] sufficient to support the fact that ”[MASK] can use Mg (2+)”.
        Query: [MASK] can use either Ca(2+) or Mg(2+) equally.
        Target: AGTR2
        Evidence: Magnesium concentrations were 2 mol/L, except for [MASK]
        (2.5 mol/L). (PMID: 11841622)
        R2E-generated Shapley value: 0.0981
        GPT4 completion: A: The evidence sentence only provides information
        about the magnesium concentration in relation to [MASK], but it does not
        provide any information about [MASK]’s ability to use either Ca(2+) or
        Mg(2+) equally. Therefore, it is <answer>IRRELEVANT</answer> with
        <confidence>HIGH</confidence>.
  4. 4.
    False negative: R2E assigns a negative Shapley value to a piece of evidence that GPT-4 labels as relevant. R2E fails to recognize that this is a relevant piece of evidence.
        Query: [MASK] is required for mitochondrial tubular network
        and cristae organization.
        Target: GHITM
        Evidence: Decreased [MASK] expression and loss of cristae organisation
        occur at an early stage of DIO and represent an IRF5-dependent mechanism
        that may contribute to loss of microenvironmental homeostasis and
        development of insulin resistance (Fig. 8 ). (PMID: 36042203)
        R2E-generated Shapley value: -0.0090
        GPT4 completion: A: The evidence sentence states that decreased [MASK]
        expression leads to loss of cristae organization, which directly
        supports the biological fact that [MASK] is required for cristae
        organization. Therefore, it is <answer>RELEVANT</answer> with
        <confidence>HIGH</confidence>.
Appendix O Further Details on Creation of Clinical Trial Outcomes Dataset
[…]
From the remaining drugs, we extracted therapeutic hypotheses, as described by a combination of a drug’s protein targets and the diseases the drug had been tested against. We discretized therapeutic hypotheses using the PharmaProjects assigned MeSH (https://www.ncbi.nlm.nih.gov/mesh/) and Entrez (Maglott et al., 2005) ontology identifiers for the genes and diseases respectively. Nelson et al. 2015 and Minikel et al. 2023 investigate the transition between all trial phases. We use only a subset that focuses on the in vivo efficacy of therapeutic hypotheses. As such, we kept only the therapeutic hypotheses related to drugs tested at Phase II or III, or pre-Registration, Registration or Launched with regulatory approval. We kept only the therapeutic hypotheses where there were no drugs in active development and therefore whose clinical efficacy could be determined.
Therapeutic hypotheses that had made it to Phase II or III and have no drugs in active clinical development were assumed to have failed to demonstrate in vivo clinical efficacy while drugs that had made it to pre-Registration and above were said to have ’succeeded’. These are the positive and negative labels in the Clinical Trial Outcomes dataset.
In constructing the Clinical Trial Outcomes dataset we made the assumption that ceased development is indicative of a therapeutic hypothesis failing to show efficacy in a human population. We highlight that there is likely to be noise in these negative labels: drug programmes can be prosecuted or abandoned for a range of commercial reasons rather than biological ones, drug programmes may fail because sponsors failed to identify an appropriate patient population, or drug programmes may fail for pharmacological reasons peculiar to the candidate molecule.
Appendix P Genetics Baseline for the Clinical Trial Outcomes Dataset
Data for the genetics baseline was downloaded from the supplementary data of Minikel et al. 2023 (https://github.com/ericminikel/genetic_support/tree/sio/data) and reproduced using the methodology described in Minikel et al. 2023, briefly summarised below.
In the supplementary data, table assoc.tsv contains the full set of genetic associations that were templated into natural language in Section 4.5.2. These already-curated genetic associations were filtered further as per Minikel et al. 2023, removing all rows with a “source” of ‘OTG‘ and an “l2g_share” $<$ 0.5.
There exists ontological mismatch between sources of genetic evidence and diseases referenced in the PharmaProjects data. As such, the Clinical Trial Outcomes dataset is joined to the genetic association data by matching exactly on gene identity, and on a measure of MeSH-MeSH similarity for diseases.
The table sim.tsv.gz contains a full list of pairwise MeSH - MeSH similarities used in this joining of datasets. The similarity measure is a composite information criterion measure of similarity on the MeSH ontology tree; see Minikel et al. 2023 for details.
The continuous score for the genetics baseline for each therapeutic hypotheses in the Clinical Trial Outcomes dataset is the maximum similarity to a genetics association across all the genetic association data, where 1 implies an exact disease-disease match and 0 means the there is no path between the entities in the MeSH ontology, or there is no genetic association data available for the gene anywhere in the genetic association data.
Appendix Q Further Details on Creation of Evidence Annotations for Clinical Trial Outcomes Dataset
We constructed Evidence Annotations for the Clinical Trial Outcomes dataset by having GPT-4 annotate (as relevant or irrelevant) all evidence for 50 Clinical Trial Outcome therapeutic hypotheses associated with trial success, as well as 50 with trial failures, both randomly sampled, resulting in 6400 query-evidence pairs with a binary annotation.
Separately and using a similar approach, we created the dataset of evidence annotations used for auditing explanations as described in Section 4.6. In this case, we computed R2E Shapley values (computed on pre-sigmoid outputs) for all retrieved evidence over all Clinical Trial Outcomes dataset examples, ordered the evidence by Shapley value, and selected the 20,000 evidence sentences with highest Shapley values. We then ran relevant/irrelevant annotations on this subset using GPT-4.
We used the combined pre-2005 literature and templated genetics corpus for both tasks. Relevant/irrelevant annotations were obtained through the use of GPT-4, using the following prompt, substituting in DISEASE_OF_INTEREST and EVIDENCE_SENTENCE:
    You are a scientific expert working on target identification in drug
    discovery.
    Your task is to use your expertise to evaluate a piece of evidence
    (referred to as EVIDENCE) for a potential drug target for a specified
    disease (referred to as DISEASE). Specifically you must indicate whether
    the EVIDENCE about a masked target (in the form of a sentence from an
    academic paper), provides relevant evidence that the drug target might be
    promising for developing a treatment for the DISEASE. If the EVIDENCE
    sentence does not make any link to the biology of the specified DISEASE,
    then it is not relevant. Please explain your reasoning first before giving
    your answer. Provide your final answer by stating either
    <answer>RELEVANT</answer> or <answer>IRRELEVANT</answer>. Please also
    indicate your confidence in your answer by writing one of
    <confidence>HIGH</confidence> or <confidence>LOW</confidence>.
    Here are some examples demonstrating proper formatting and reasoning in
    a response:
    <example>
    H: Your DISEASE of interest is Sarcopenia.
    Here is the EVIDENCE sentence, containing a masked target, to assess:
    <evidence>
    Many studies also described exercise-induced increases in transcriptional
    and translational levels of FGFR1, [MASK], and/or KLB [29,33,35,36].
    </evidence>
    A:The evidence sentence makes no indication that [MASK] plays a role in
    Sarcopenia, therefore it is <answer>IRRELEVANT</answer> with
    <confidence>HIGH</confidence>.
    </example>
    <example>
    H: Your DISEASE of interest is Amyotrophic Lateral Sclerosis.
    Here is the EVIDENCE sentence, containing a masked target, to assess:
    <evidence>
    Therefore, further study is needed to clarify where [MASK] functions
    during lysosome trafficking and neurite outgrowth.
    </evidence>
    A:The evidence sentence implies that [MASK] may play a role in biological
    mechanisms directly related to ALS, but the phrasing is ambiguous, therefore
    it is <answer>RELEVANT</answer> with <confidence>LOW</confidence>.
    </example>
    <example>
    H: Your DISEASE of interest is Lung Adenocarcinoma.
    Here is the EVIDENCE sentence, containing a masked target, to assess:
    <evidence>
    Conversely, [MASK] mRNA and protein expression is higher in a variety of
    tumor tissues, including lung cancer [ 7 , 20 ].
    </evidence>
    A:The evidence sentence states that [MASK] expression is upregulated in tumor
    tissues including our specific cancer type of lung cancer, therefore it is
    <answer>RELEVANT</answer> with <confidence>HIGH</confidence>.
    </example>
    Now here is your real task.
    Your DISEASE of interest is {DISEASE_OF_INTEREST}.
    Here is the EVIDENCE sentence, containing a masked target, to assess:
    <evidence>
    {EVIDENCE_SENTENCE}
    </evidence>
Appendix R Examples of Evidence Annotations for Clinical Trial Outcomes Dataset
Here we show an example of each of true positive, true negative, false positive, and false negative R2E evidence Shapley values versus GPT-4 relevant/irrelevant Evidence Annotations, for Clinical Trial Outcomes data.
  1. 1.
    True positive: R2E assigns a large, positive Shapley value to a piece of evidence that GPT-4 labels as relevant. R2E and GPT-4 recognise this as a relevant piece of evidence.
        Query: [MASK] is a promising drug target for the treatment of
        myelodysplastic syndromes.
        Target: BIRC2
        Evidence: [MASK] protein expression was found to be present in a majority
        of leukemic blasts from untreated patients, but absent in normal
        differentiating myeloid progenitor cells. (PMID: 14666661)
        R2E-generated Shapley value: 0.4057
        GPT4 completion: A: The evidence sentence indicates that [MASK] protein is
        expressed in leukemic blasts, which are abnormal cells associated with
        leukemia, a type of blood cancer. Myelodysplastic Syndromes (MDS) are a
        group of disorders caused by poorly formed blood cells or ones that don’t
        work properly, which can often progress to leukemia. The fact that [MASK]
        is not present in normal myeloid progenitor cells, which are cells that
        give rise to all types of blood cells, suggests that it could be involved
        in the abnormal development of these cells in MDS. Therefore, it is
        <answer>RELEVANT</answer> with <confidence>HIGH</confidence>.
  2. 2.
    True negative: R2E assigns a negative Shapley value to a piece of evidence that GPT-4 labels as irrelevant. R2E and GPT-4 recognise this as an irrelevant piece of evidence.
        Query: [MASK] is a promising drug target for the treatment of crohn
        disease.
        Target: LANCL2
        Evidence: Because of the structural similarity to LanC, we postulate that
        [MASK] may play a role as a component of a peptide-modifying complex.
        (PMID: 11762191)
        R2E-generated Shapley value: -0.0160
        GPT4 completion: A: The evidence sentence suggests that [MASK] may be
        involved in peptide modification, but it does not provide any direct link
        or relevance to Crohn Disease. Therefore, it is <answer>IRRELEVANT</answer>
        with <confidence>HIGH</confidence>.
  3. 3.
    False positive: R2E assigns a large, positive Shapley value to a piece of evidence that GPT-4 labels as irrelevant. R2E seems to misunderstand an abbreviation that GPT-4 is able to correctly parse.
        Query: [MASK] is a promising drug target for the treatment of chronic
        myelomonocytic leukemia.
        Target: MAP3K1
        Evidence: This observation indicates that [MASK] may be a useful
        therapeutic target in the treatment of CML. (PMID: 14586403)
        R2E-generated Shapley value: 0.2329
        GPT4 completion: A:The evidence sentence suggests that [MASK] could
        be a potential therapeutic target for Chronic Myeloid Leukemia (CML).
        However, the disease of interest is Chronic Myelomonocytic Leukemia,
        not CML. Despite both being types of leukemia, they are distinct diseases
        with different pathologies. Therefore, the evidence is
        <answer>IRRELEVANT</answer> with <confidence>HIGH</confidence>.
  4. 4.
    False negative: R2E assigns a negative Shapley value to a piece of evidence that GPT-4 labels as relevant. R2E fails to recognize that this is a relevant piece of evidence.
        Query: [MASK] is a promising drug target for the treatment of heart
        failure.
        Target: RXFP1
        Evidence: However, a recent study has identified the relaxin receptor as a
        G-protein-coupled orphan receptor, [MASK], that can couple to adenylate
        cyclase to increase cAMP ( Hsu et al ., 2002 ). (PMID: 12381685)
        R2E-generated Shapley value: -0.0167
        GPT4 completion: A: The evidence sentence suggests that [MASK] is a
        G-protein-coupled orphan receptor that can couple to adenylate cyclase
        to increase cAMP. This is relevant to heart failure as G-protein-coupled
        receptors and cAMP signaling pathways are known to play crucial roles in
        cardiac function and heart failure pathology. Therefore, it is
        <answer>RELEVANT</answer> with <confidence>HIGH</confidence>.
Appendix S Details on Cleaning MeSH Terms For Templating
The canonical MeSH name for each MeSH identifier is used in templates for generating R2E queries in the Clinical Trial Outcomes evaluation and generating genetics-derived sentences for augmenting the literature evidence.
The MeSH names underwent light reformatting to align them to how they might be expressed in natural language. Simply, we lowercase the MeSH name, split on commas, and reverse order the resulting list. So for example the MeSH name Leukemia, Myelomonocytic, Chronic becomes chronic myelomonocytic leukemia.
Appendix T Further Results on Predicting Clinical Trial Outcomes and Genetic Evidence
Table S8 shows that the best performing R2E model retrieving from literature alone, was R2E with bias correction and without temporal training (see Appendix K). As such, in Table 3, results without temporal training and with bias correction are shown.
Table S9 shows that the trends in AUROC are similar if we restrict the post-2005 therapeutic hypotheses as per Minikel et al. 2023. We restrict therapeutic hypotheses to only those diseases with ’genetic insight’. Diseases were only kept if there was at least one genetic association between a gene and a disease with a similarity of $>$ 0.7, leaving 4,226 therapeutic hypotheses.
The genetics baseline has a widely published Relative Success (RS) of approximately 2 in predicting Clinical Trial success. This RS estimate is derived from a set of high confidence predictions (500 positive predictions across the 4,226 therapeutic hypotheses) from the genetics baseline (using a similarity threshold of $>$0.8 as per Minikel et al. 2023). R2E-cor predicting on literature obtained an RS equivalent to the genetics baseline (1.98; 95% CI (1.78, 2.21)), at a score threshold of 0.6, but with almost twice as many positive predictions (991 positive predictions across the 4,226 therapeutic hypotheses).
When calculating AUROC, the genetics baseline ranks every disease-gene pair in the Clinical Trial Outcomes data, based on the similarity of the trial disease to the nearest (most similar) genetic baseline disease, rather than binarising above and below a similarity of 0.8. The combination of low AUROC but high RS can be explained by the genetics baseline performing badly against the majority of therapeutic hypotheses for which there is no direct genetics evidence (i.e. the similarity of the nearest disease with genetic evidence is $<$ 0.8).
Table S10 shows that in the multi-modal context (with a corpus of sentences from the biomedical literature and from the genetics data), forcing retrieval of genetics evidence does not change the AUROC.
Figure S4 shows that there is substantial variation in performance across disease areas and modality. We highlight that the magnitude of difference in performance between R2E retrieving from genetics alone and R2E retrieving from literature, varies by disease area. This may indicate disease areas for which alternative modalities could be valuable to template into text.
Figure S5 shows the distribution of MeSH-MeSH ontological similarity, as calculated by Minikel et al. 2023, for query-evidence pairs with high Shapley scores (see Section 4.6) and where genetics evidence was annotated as relevant by GPT-4 (n=527) . Note that when calculating RS in Minikel et al. 2023, the threshold similarity for linking between therapeutic hypotheses and genetic association data is 0.8. Some examples of the “soft” semantic matching between the query and the genetics evidence for MeSH pairs with a similarity $<$ 0.2 are: (erythrocyte count and anemia), (eosinophilia and asthma), (astrocytoma and brain neoplasms). This highlights the shortcomings of a universal threshold based on ontological similarity metrics versus semantic matching and reasoning through natural language.
Table S8: Temporal training and Clinical Trial Outcomes: AUROC for R2E-Uncorrected and R2E-Corrected with and without enforcement of retrieval of literature evidence pre-query during training. Retrieval at inference time remains the same (i.e. pre-2005 evidence).
Model Train-time retrieval
Standard Pre-query
R2E-uncor 0.634 0.626
R2E-cor 0.639 0.628
Table S9: Clinical Trial Outcomes on therapeutic hypotheses with genetic insight: AUROC for R2E retrieving from literature-alone, genetics-alone, or both; in comparison to baselines, when subsetting therapeutic hypotheses just to those where the disease has at least one genetic association in the genetics baseline.
Model Modality AUROC
Genetic Genetics 0.588
Freq Literature 0.552
MCS Literature 0.634
MLM Literature 0.638
R2E-cor Literature 0.643
R2E-uncor Genetics 0.618
R2E-cor Combined 0.647
Table S10: Methods of multimodality for Clinical Trial Outcomes: AUROC for R2E-Uncorrected and Corrected with three different methods of multi-modalility: (1) Retrieve from a single corpus containing both genetics and literature sentences (single index); (2) Retrieve up to four sentences from the genetics corpus - where possible - and retrieve the remaining sentences from the literature corpus (separate index); and (3) R2E scores evidence from the genetics and the literature corpora separately and the final score is the mean of the two (post-hoc aggregation).
Model Modality Method AUROC
R2E-uncor Combined Single index 0.631
R2E-cor Combined Single index 0.633
R2E-uncor Combined Seperate index 0.631
R2E-cor Combined Seperate index 0.633
R2E Combined Post-hoc aggregation 0.633
Figure S4: R2E performance across disease areas. AUROC in each PharmaProjects annotated disease area, with R2E retrieving from literature-alone (R2E-cor), genetics-alone (R2E-uncor), or both (R2E-cor); in comparison to the genetic baseline. The number of successes over total number of therapeutic hypotheses for each disease area are given in brackets.
Figure S5: MeSH-MeSH ontological similarity of query-genetics evidence pairs with high Shapley scores. Histogram showing the distribution of similarities between MeSH terms for diseases in genetics evidence annotated as “Relevant” by GPT-4, and diseases in the clinical trial therapeutic hypothesis queries, when the evidence had a high Shapley score (see Section 4.6). The red dashed line indicates the threshold used in the Minikel et al. 2023 genetics baseline to assign genetic evidence to a therapeutic hypothesis.
Appendix U Examples of Auditing of Evidence for Clinical Trial Outcomes Dataset
Here we show examples of query-evidence pairs that GPT-4 annotated as irrelevant and to which R2E assigned a large positive Shapley value, as identified during the auditing experiments detailed in Section 4.6. With each example we also report the overall R2E prediction score for the associated query-target pair before and after the auditing process. The change in score from before to after the auditing process accounts for the removal of the specified “false positive” (according to GPT-4) piece of evidence, as well as any other query-evidence “false positive” pairs removed by the audit. We also include the final clinical trial status of the associated query-target pair in each example.
  1. 1.
    Negative evidence example: GPT-4 correctly identifies the negation in the evidence sentence, which suggests that there is not good evidence for the involvement of [MASK] in Tourette’s syndrome. R2E fails in this case and assigns a large positive Shapley value to this irrelevant piece of evidence.
        Query: [MASK] is a promising drug target for the treatment of tourette
        syndrome.
        Target: HTR1A
        Clinical Trial Outcome: Failure
        R2E score pre-audit: 0.9831
        R2E score post-audit: 0.9395
        Evidence: In conclusion, our study does not provide evidence that the
        [MASK] gene plays either a major or a minor role in the genetic
        predisposition to schizophrenia, bipolar affective disorder, or Tourette’s
        syndrome. (PMID: 8546152)
        R2E-generated Shapley value: 0.0371
        GPT4 completion: A:The evidence sentence explicitly states that the [MASK]
        gene does not play a major or minor role in the genetic predisposition to
        Tourette’s syndrome, therefore it is <answer>IRRELEVANT</answer> with
        <confidence>HIGH</confidence>.
  2. 2.
    Failed entity linking example: R2E incorrectly understands ”ICH” to refer to the disease of interest ”ichytosis”, which is a skin disorder, and therefore assigns a large Shapley value to this irrelevant piece of evidence. GPT-4 correctly infers that ”ICH” refers here to intracerebral hemorrhage based on the surrounding context in the evidence sentence, and is therefore irrelevant to ichytosis.
        Query: [MASK] is a promising drug target for the treatment of ichthyosis.
        Target: TNF
        Clinical Trial Outcome: Success
        R2E score pre-audit: 0.8920
        R2E score post-audit: 0.7040
        Evidence: [MASK] likely plays a central and pathogenic role during ICH,
        because antisense oligonucleotides that target [MASK] reduced [MASK]
        protein levels and neuronal cell death within the hematoma as well as
        neurobehavioural deficits following ICH. (PMID: 11409424)
        R2E-generated Shapley value: 0.1088
        GPT4 completion: A: The evidence sentence suggests that [MASK] plays a
        central and pathogenic role during ICH, and that reducing [MASK] protein
        levels can alleviate symptoms. However, the ICH referred to in the
        sentence is likely intracerebral hemorrhage, not ichthyosis. Therefore,
        the evidence is <answer>IRRELEVANT</answer> with
        <confidence>HIGH</confidence>.
  3. 3.
    Incorrect understanding example: R2E assigns a large positive Shapley value to a piece of evidence that seems to have no relevance (perhaps reasoning incorrectly based on the shared ”retin” token in the disease name and the compound name). GPT-4 correctly identifies the evidence as irrelevant.
        Query: [MASK] is a promising drug target for the treatment of retinitis
        pigmentosa.
        Target: NGFR
        Clinical Trial Outcome: Failure
        R2E score pre-audit: 0.6419
        R2E score post-audit: 0.3103
        Evidence: [MASK] was solubilized in methylene chloride and fluorinated
        using diethylaminosulfurtrifluoride (DAST) as the fluorinating agent
        (Fig. 1) asdescribed for the synthesis of all- trans-retinoyl-fluoride
        [25]. (PMID: 11762191)
        R2E-generated Shapley value: 0.3807
        GPT4 completion: A: The evidence sentence describes a chemical process
        involving [MASK], but does not link this process to the biology or
        pathology of Retinitis Pigmentosa. Therefore, it is
        <answer>IRRELEVANT</answer> with <confidence>HIGH</confidence>.
  4. 4.
    Indirect evidence example: GPT-4 is arguably too restrictive in its definition of relevance here. B-cell chronic lymphocytic leukemia (B-CLL) and mantle-cell lymphoma share enough significant disease biology that this is still a potentially relevant piece of evidence. This bias against indirect evidence was an issue we saw in several auditing examples we examined, and could be mitigated by adjusting the prompting strategy to GPT-4 in future work.
        Query: [MASK] is a promising drug target for the treatment of mantle-cell
        lymphoma.
        Target: TNFRSF10B
        Clinical Trial Outcome: Failure
        R2E score pre-audit: 0.5400
        R2E score post-audit: 0.0381
        Evidence: Cells from 10 B-CLL patients previously studied by flow
        cytometry and from an additional three B-CLL patients (L-244, L-243 and
        L-249) were also positive for [MASK] protein by Western blot analysis
        ( Figure 3 ). (PMID: 11753607)
        R2E-generated Shapley value: 0.5914
        GPT4 completion: A: The evidence sentence indicates that [MASK] protein is
        present in B-CLL (B-cell Chronic Lymphocytic Leukemia) patients. However,
        B-CLL is a different type of lymphoma than Mantle-Cell Lymphoma. While they
        are both types of B-cell lymphomas, they have different characteristics and
        may not share the same drug targets. Therefore, the evidence does not
        directly link [MASK] to Mantle-Cell Lymphoma. It is
        <answer>IRRELEVANT</answer> with <confidence>HIGH</confidence>.